OpenStudy (anonymous):

Which is the equation of a circle with center (–2, 3) and radius r = 5? (x + 2)² + (y – 3)² = 10 (x + 2)² + (y – 3)² = 25 (x – 2)² + (y + 3)² = 10 (x – 2)² + (y + 3)² = 25

OpenStudy (anonymous):

The equation of a circle with center (a,b) and radius r is (x-a)^2 + (y-b)^2 = r^2.

OpenStudy (anonymous):

(x + 2)² + (y – 3)² = 25

OpenStudy (anonymous):

the \[r^{2}\] at the end